Massachusetts Institute of Technology Handout ?? 6.042J/18.062J: Mathematics for Computer Science February 13, 1997 Professor David R. Karger

Office hours, etc.

1 Recitations

Recitation Time Room TA
1a  1pm 26-210 Daniele
1b  1pm 26-302 Alex
2a  2pm 26-210 Daniele
2b  2pm 34-301 Alex
3a  3pm 34-302 Daniele
3b  3pm 34-301 Alex

2 Tutorials

Tutorial Time Room TA
1A  1pm 2-131 Dimitri
1B  1pm 2-135 Terri
1C  1pm 2-139 Dave
2A  2pm 34-302 Dimitri
2B  2pm 34-303 Terri
3A  3pm 34-303 Terri
3B  3pm 34-304 Dave
3C  3pm 34-302 Dimitri
4  ?  ?  Dave

3 Office hours

Alex  Friday 4-6pm  NE43-308
Daniel  Monday 10am-Noon NE43-342
David, Terri Monday 4-6pm  NE43-3rd floor lounge (outside NE43-322)
Dimitri  Monday 5-7pm  NE43-3rd floor lounge (outside NE43-322)